OpenStudy (anonymous):

Check my answer? Given the proportion a/b = 8/15, which ratio completes the equivalent proportion a/8 = ? My answer: 15/b @KirbyLegs

OpenStudy (anonymous):

@dc104

OpenStudy (anonymous):

This is another cross multiplication XD 8/15 = a/8 15a = 64 a = 64/15 They are not trying to make you find b. They only want you to find the proportion of the division terms given.

OpenStudy (anonymous):

The answer choices they gave me were 15/b, b/15, 8/15, and a/15.

OpenStudy (anonymous):

@KirbyLegs has given you a, they want a/8

OpenStudy (anonymous):

a = 64/15, so a/8 = (64/8)/15 = 8/15